OpenStudy (anonymous):

1. Sin ( B-A) is equal to _____, when B = 270 degrees and A is an acute angle.

jimthompson5910 (jim_thompson5910):

hint: sin(B - A) = sin(B)*cos(A) - cos(B)*sin(A)

OpenStudy (anonymous):

whta's the answer?

OpenStudy (anonymous):

Csc 520 is equal to

OpenStudy (anonymous):

\[\sin \left( B-A \right)=\sin \left( 180+90-A \right)=-\sin \left( 90-A \right)=-\cos A\]

OpenStudy (anonymous):

\[\csc 520=\csc \left( 360+160 \right)=\csc 160=\csc \left( 90+70 \right)=\sec 70\]